Scores on a test are normally distributed with a mean of 78 and a standard deviation of 6. A student is selected at random. Which has the greatest probability?


Select one:
a.
The student's score is greater that 89.

b.
The student's score is less than 69.

c.
The student's score is between 63 and 72.
d.
The student's score is between 78 and 81.

Answer by Theo(13342)

mean is 78.
standard deviation is 6.

a.)
score greater than 89 has z-score of (89 - 78) / 6 = 11/6 = 1.83.
area to the right of that is equal to .0336



b.)
score less than 69 has z-score of (69 - 78) / 6 = -9/6 = -1.5
area to the left of that is equal to .0668



c.)
low z-score = (63 - 78) / 6 = -2.5
area to the left of that is equal to .00621
high z-score = (72 - 78) / 6 = -1
area to the left of that is equal to .15866
area in between is equal to .1524



d.)
low z-score = (78 - 78) / 6 = 0
area to the left of that is equal to .5
high z-score = (81 - 78) / 6 = .5
area to the left of that is equal to .69146
area in between is equal to .1915